# LPSim: Large Scale Multi-GPU Parallel Computing based Regional Scale Traffic Simulation Framework

### v1

@@ 1 Byte in Memory = 1 Meter in Road Segments
Input: $\mathcal{R}$ (Road Status), $t_{depart}$ (Departure Time), $\mathcal{V}$ (Vehicle Type), $e_{current}$ (Current Edge), $v$ (Speed), $\mathcal{I}$ (Intersection Status)
@@
5:         Start moving
6:         if New edge detected then
@@
10:              $d_{front}\leftarrow 2\cdot\Delta t\cdot v$
@@
18:              Evaluate lane change
@@
20:                  if Gap acceptable then
@@
22:return $\mathcal{P},v,a$
3.2 Architecture of single GPU-based traffic simulation
3.2.1 GPU Kernel Design for Vehicle and Traffic Management
In detail, as shown in the Vehicle Propagation Algorithm 1, the $\mathcal{R}$ Road Status array represents the occupancy status or speed of vehicles on the road, while $\mathcal{I}$ (Intersection Status) denotes the graph nodes. This paper employs an adjacency list to map the directed connectivity of nodes, with list entries pointing to arrays that describe road segments (e.g., a 1D array for a 4-lane, 8-meter road segment will be the dimension of 1 X (4*8), resulting in dimensions of 1 X 32 as shown in Fig 2.) using Char Type in C++ to indicate road status (occupied, empty, and vehicle speed, etc.) , the real road segment is mapped into a 1D array organizing lanes in order. The process begins with each thread checking the road status and vehicle’s departure time to decide whether to depart or wait. Once the vehicle is active, the thread orchestrates the vehicle’s progression, revising its position and velocity in response to the immediate road conditions, such as the presence and distance of preceding vehicles. Additionally, the thread evaluates the vehicle’s interaction with intersections and potential lane changes.
@@ 3.3 Design of multiple GPU-based simulation framework
The propagation of vehicles and the complete road network across multiple GPUs calls for inter-GPU communications to synchronize state and share data. When a vehicle is not located in the ghost zone, the simulator checks whether its next node belongs to the ghost zone. If not, the simulation continues to a single-GPU scenario, eliminating the need for cross-GPU computations. Conversely, if the vehicle is about to enter the ghost zone, it is duplicated in the ghost zone of the destination GPU. If a vehicle is already in a ghost zone, it is assured that its next node is out of the ghost zone which means after the current edge traversal, the vehicle will be on the next GPU and leave the original GPU. At this juncture, it is determined whether the node resides within the current GPU’s domain. If it does, the situation reverts to a single-GPU model as described in Session 3.2 again. If the node is outside the current GPU, the vehicle is marked for removal, and subsequently eliminated post the completion of the simulation timestep. The whole procedure is summarized in Figure 3, intersections within the road network are represented by circles, and we use an adjacency list to illustrate the directed relationships between these intersections. The entries in the list link to the arrays that define road segments. For example, a 4-lane 8-meter road segment is depicted as a 1D matrix with a dimension of 1 x 32 as shown in Fig. 2, showing how real road segments, represented by lines in Fig. 3, are systematically translated into 1D matrices to sequentially organize lanes. In the process of selecting ghost zones, we replicate the entire edge (represented as a 1D array) across both GPUs. To ensure exclusive byte allocation to each vehicle, we transfer vehicles one at a time from one edge to another if the first-byte memory of the downstream edge is not occupied.
@@ 3.3.1 Graph Partitioning
In large-scale multi-GPU parallel simulations, it is essential to distribute graph data across multiple GPUs to share computational resources. We aim for an efficient allocation of extensive graph data, managing computational resources while minimizing inter-GPU communication. In this part, we first state our problem and propose an approximate optimization formulation of the graph partitioning task. Since it is difficult to solve for our network with $200k$ node size, we introduce two applicable graph partitioning strategies to solve the problem. The choice of two graph partitioning methods is closely related to the available computational resources of the single GPU, and the inherent approximation idea of the two methods comes from different perspectives. The framework is summarized as shown in Figure 4, which illustrates that when we have insufficient computational resources, which means that we are creating many more threads than CUDA cores, it leads to a computation-bound scenario. We propose using balanced partitioning to allocate computation evenly across GPUs. On the contrary, with ample computational power, the challenge shifts to minimizing communication overhead, necessitating unbalanced partitioning. Our framework consists of three parts: identifying the difficulties in the partition problem, simplifying it to an optimization formulation, and using efficient algorithms to solve the approximation problem. The left side of Fig. 4 is about modeling the problem, the right side is about how to formulate the problem and solve it, and the middle part is about the detailed components of our solutions.
@@
    Outlier Detection: For those cities never visited in the past time period, we allocate them to the nearest subgraph.
We present an example of partition results for two distinct time periods, showing both balanced and unbalanced graph partitionings in Figures 5 and 6, respectively. These partitions are aligned with the dynamics of the real world traffic flow. Specifically, areas such as the Bay Bridge and Treasure Island are unified within a single cluster due to the incessant flow of traffic in both balanced and unbalanced partition cases because of their heavy traffic throughout the day, underscoring the impracticality of division. In the balanced partition scenario, the map is divided into northern and southern segments by the Bay Bridge area, delineating a clear geographic split. On the other hand, the unbalanced partition strategy, particularly with 4 and 8 clusters, mirrors the Bay Area’s administrative boundaries more closely. For the partition featuring four clusters, the blue, green, and orange parts represent North Bay, East Bay, and South Bay, respectively, while the red segment encompasses the San Francisco and Peninsula areas, along with portions of North and East Bay, interconnected by the Golden Gate Bridge and the Bay Bridge. In the scenario with eight clusters, each segment closely approximates the distinct counties within the Bay Area, with the exception that San Francisco and Marin counties are in one segment.
@@ 3.3.2 Inter-GPU Communication
The additional cost incurred in a multiple-GPU 7 setup, compared to a single-GPU system, arises predominantly from the need for communication between GPUs. We facilitate communication between multiple GPUs to handle two types of data: vehicles and road networks within ghost zones. Vehicle data transmission occurs when a vehicle enters a ghost zone, while road networks data communication is necessary to maintain consistency across different GPUs when a vehicle moves within a ghost zone. While the use of arrays to store vehicle data, managed via cudaFree and cudaMalloc, presented an intuitive approach, it revealed significant drawbacks, especially in handling variable data sizes. As illustrated in Scenario 1 of Figure 8, allocating excessive memory leads to inefficiency, whereas insufficient allocation, shown in Scenario 2, poses the risk of array index out of bounds. Moreover, this method was hampered by the time-consuming processes involved in frequent memory allocations and deallocations. To efficiently manage the variable-length vehicle data, which needs to be transferred between GPUs, we employ a more effective approach, utilizing device_vector from the Thrust library, the CUDA C++ template library. Thrust’s device_vector is designed for GPU contexts and allows dynamic resizing of its contained elements. This is achieved through efficient memory management strategies, which minimize the overhead of reallocating memory when the device_vector size changes.
@@
Building on the detailed methodologies for vehicle replication and deletion, and road network data consistency, we’ve established a robust system for managing GPU communication. The following section demonstrates the relatively minimal time impact of communication operations compared to read and write processes from one GPU to its local memory, underscoring the practicality and efficiency of our multi-GPU approach. For read and write operations, the cost was assessed by reading and writing large arrays, where each thread manipulates a single array value. In contrast, communication operations were evaluated by replicating a long array from one GPU to another, using both Device-Host-Device (D-H-D) and Peer-to-Peer (P2P) methods. P2P, especially via NVLink, is claimed faster than D-H-D, though its availability depends on the machine’s hardware architecture. Experimental results on V100 GPUs equipped with NVLink revealed that the time cost for P2P communication is approximately threefold that of read-write operations, shown as Figure 9, indicating a reasonably efficient system. In the upcoming performance section, we will elaborate using the example of San Francisco Bay Area, how graph partitioning techniques and other strategies effectively minimize communication overhead to a mere fraction (about 1‰) of the total computational load. In summary, while inter-GPU communication does introduce additional time cost, this is relatively minimal compared to the extensive read and write operations, thereby justifying the use of multiple GPUs for their significant performance benefits.
@@
Figure 10: The Method of Vehicle Replication and Deletion
3.4 Scalability Theory of Traffic Simulation
In this section, we propose a way to combine Bulk synchronous parallel (BSP) and normalized runtime across different GPU hardware to measure the scalibility on single-GPU machine. And we also introduced the implications from Amdahl’s law for quantifying the scalibility of multi-GPU traffic simulation
3.4.1 BSP modeling for single-GPU approach performance benchmarking
Currently, the graph computing frameworks largely follow the Bulk Synchronous Parallel (BSP) computation model as shown in figure 11. This model divides computations into a series of iterations called supersteps. The cost of a superstep is the sum of the following three terms:
  • 1.
    Cost of the longest-running local computation
  • 2.
    Cost of global communication between the processors
  • 3.
    Cost of the barrier synchronization at the end of the superstep.
We use the following notations to show the overall computation cost of the BSP model.
  • 1.
    $w_{i}$ time spent for the local computation in process $i$
  • 2.
    $h_{i}$ number of messages sent or received by process $i$
  • 3.
    $l$ time spent for synchronization
The cost of a whole BSP algorithm is the sum of the cost of each superstep: $W+Hg+Sl=\sum_{s=1}^{S}w_{s}+g\sum_{s=1}^{S}h_{s}+Sl$, where $S$ is the number of supersteps.
Figure 11: BSP model Cheatham et al. [1996]
3.4.2 Normalizing runtime across different GPU hardware
After introducing the computation model, we consider the normalization of runtime across different GPU hardware, which refers to measuring the runtime in a way that is as independent of hardware specifications as possible. We argue that if the runtime is broken down following the BSP model, it can roughly be normalized using key hardware features. Specifically, we normalize the computation time with the clock rate Bakhoda et al. [2009], the communication time with the bandwidth Pearson et al. [2019], and the synchronization time with bandwidth Zhang et al. [2020]. We will give a brief description below.
For the normalized runtime $\bar{t}_{\mathrm{run}}$, it is given by
$\bar{t}{\mathrm{run}}=\frac{t{\mathrm{comp}}}{R_{1}/R_{i}}+\frac{t_{\mathrm{comm}}}{B_{1}/B_{i}}+\frac{t_{\mathrm{sync}}}{R_{1}/R_{i}}$ (2)
Where $R_{i}$ refers to the clock rate in GHz for machine $i$, where $R_{1}$ is the clock rate of the benchmark machine (machine with the lowest number of CUDA cores), and $B_{i}$ refers to the bandwidth in GB/s for the machine $i$, where $B_{1}$ is the bandwidth of the benchmark machine. The number of cores, bandwidth, and the clock rate are independent. The purpose of Equation 2 is to enable a fair comparison across different runs that vary in the number of cores, bandwidth, and clock rates. By normalizing for clock rate and bandwidth, it isolates the measurement of speedup to focus solely on the influence of core count.
3.4.3 An application of Amdahl’s law for multi-GPU scalibility benchmarking
After introducing the single machine computation model, we move to the parallel computing analysis theory in this part. The main tool we want to use is the Amdahl’s law Hill and Marty [2008]. It claims that the theoretical speedup $S_{p}$ that can be achieved with $p$ processors parallelization is
$S_{p}=\frac{1}{f+\frac{1-f}{p}}$ (3)
where $f$ refers to the fraction of the computation that must be executed sequentially (i.e., cannot be parallelized), and $1-f$ is the fraction that can be executed in parallel. We note that when $f=0$, $S_{p}=p$ and when $f=1$, $S_{p}=1$.
Another interpretation of the equation (3) is $S_{p}=\frac{t(s,1)}{t(s,p)}$. This equation represents the actual speedup achieved in practice. Here, $t(s,1)$ denotes the sequential runtime of the computation and $t(s,p)$ represents the runtime with $p$ processors.
The equation above shows that as the number of processors increases, the speedup $S_{p}$ of the computation will eventually level off due to the sequential component $f$, which cannot be further parallelized. Therefore, Amdahl’s law suggests that achieving high speedup requires identifying and optimizing the parallelizable parts of a computation while minimizing the sequential components.
To accurately evaluate the performance of a parallel processing system, it is essential to measure the runtime of the computation as a function of the number of processors employed. In the context of parallel algorithms, the evaluation of system performance is a multifaceted task that requires considering various factors such as scalability, communication overhead, load balancing, and problem size. One commonly used approach to assess the efficacy of parallel processing is to compare the run time of the computation for different numbers of processors DeLee [2018], specifically the sequential runtime and the parallel runtime. Sequential runtime denotes the time it takes to execute the portion of the algorithm that cannot be parallelized, while parallel runtime represents the time it takes to execute the portion of the algorithm that can be parallelized.
3.5 Instruction Roofline Model
At the end of this part, we introduce the necessity of the Instruction Roofline Model Ding and Williams [2019], which is an evolution of the Roofline model, to uncover performance bottlenecks and enhance execution efficiency. The Roofline model offers a comprehensive framework for evaluating computing performance by integrating floating point capabilities, arithmetic intensity, and memory bandwidth into a two-dimensional representation. It excels in distinguishing whether a computing task is memory-bound or compute-bound. Ding et al. Ding and Williams [2019] proposed the Instruction Roofline Model for GPUs. Compared with the traditional roofline model, this model shifts the metric to instruction counting, enabling the capture of the mix of floating-point, integer instructions, instruction throughput, and thread predication.
To show the specifics of the Instruction Roofline Model, we draw the picture, in which the warp Instructions per Sector is the x-axis and warp billions of instructions per second (GIPS) is the y-axis. Memory access is coalesced into a sector, which is 32 bytes in size. A warp-level load may trigger between one to 32 sector transactions based on the memory access patterns, making the ”sector” a natural unit for memory access analysis. In the Instruction Roofline Model, the roofline consists of two distinct parts: a horizontal line and sloping lines. The horizontal line embodies the peak performance, which is the maximum warp GIPS achievable by the machine. The sloping lines, on the other hand, reflect the memory bandwidth. The Instruction Roofline Model is described in
$GIPS\leq\min\left\{\begin{array}[]{l}\text{ Peak GIPS }\\ \text{ Peak GTXN/s }\times\text{ Instruction Intensity }\end{array}\right.$
where GTXN/s is billions of transactions per second.
For optimal GPU performance within the Roofline Model framework, a program must effectively parallelize by adhering to GPU-specific principles: it should capitalize on data parallelism, optimize memory access for coalescence, reduce warp divergence, enhance arithmetic intensity to favor compute over memory operations, and strategically utilize the GPU’s memory hierarchy to approach the peak GIPS ceiling. These practices ensure that a GPU can perform at its maximum potential, staying within the compute-bound region of the Roofline, indicative of good performance.
@@ 4.1 Experimental Data Source
Our analysis focused on a comprehensive dataset encompassing over 19.4 million recorded trips, which is a typical weekday with no significant events or seasonal impactsChan et al. [2023]. Each entry in this dataset was categorized by origin, destination, mode of transportation, and several other critical attributes. A significant portion of our research was devoted to examining car trips, specifically categorized as Single-Occupancy Vehicle (SOV), High Occupancy Vehicle with two occupants (HOV 2), and High Occupancy Vehicle with three or more occupants (HOV 3+). Upon applying these criteria, our scope narrowed down to 15.6 million car trips, constituting 80.4% of the total dataset. This subset provided a substantial basis for analyzing vehicular movement patterns within the region, offering insights into the predominant vechicle propagation and its implications for urban planning and traffic management.
@@ 4.2 Experimental Results
1) Normalized performance with single-GPU machines having different number of CUDA cores. For this part of the experiment, due to the limitations of GPU memory, the SFCTA demand Bent et al. [2010] cannot be fully processed on a single GPU, and as evidenced in Table 7, careless graph partitioning will also lead to failure. Consequently, we used data from the Bay Area Metropolitan Transportation Commission (MTC) Cox [2013] to model origin-destination demand, specifically concentrating on automobile trips during the morning hours from 5 AM to 12 PM. This includes private vehicles, Transportation Network Companies (TNCs) , and drives to transit stops, excluding public transit and large freight trucks. Although this approach does not perfectly mirror real-world congestion—omitting around 250K trips—it still provides a broad overview with about 3 million trips considered for demonstrating the scalibility of the model .
Table 5: Related Machine Specifications
GPU Name Tesla M60 Tesla T4 Tesla V100-SXM2
Number of Streaming multiprocessors (SMs) 8 40 80
Core Clock (GHz) 1.18 1.59 1.53
CUDA Cores/SM 256 64 64
Bandwidth (GB/s) 113 215 835
Memory Amount $8\mathrm{~{}GB}$ 16 GB 32 GB
GPU architecture Maxwell Turing Volta
Special Features dynamic parallelism
real-time ray tracing
technology
high-speed communication
between GPUs and improved memory bandwidth
Table 6: Normalized runtime on single GPUs
GPU Number of Cores Time on Communication (s) Time on Computation (s) Time on Synchronization (s) Total Runtime (s) Normalized Runtime (s)
Tesla V100-SXM2 5120 36.2192 48.0529 0.7506 85.0227 335.4897992
Tesla T4 2560 139.757 144.854 3.782 288.393 465.9977431
Tesla M60 2048 267.301 377.214 4.405 648.92 648.92
The specifications of the machines used for this study are shown in Table 5. As demonstrated in Figure 12 whose detailed stats is shown in Table 6, the simulation shows a positive correlation between the number of processors used and its overall speedup. However, it should be noted that this increase in speedup is not sustained at the same rate as the initial gains observed with a smaller number of processors Hill and Marty [2008]. As discussed during our discussion in Session 3.4.2, we established that the normalized runtime offers a balanced comparison, assuming that computation, communication, and synchronization times are directly tied to clock rate and bandwidth. However, this analysis did not take into account the architectural variations across different GPUs. This oversight leads to the observation that the Tesla T4 achieves a speedup surpassing the ideal. Despite this, the overarching conclusion indicates that our model benefits from a reduction in execution time as the number of GPUs scales up.
Figure 12: Log-Log Plot of GPU Performance: CUDA Cores vs. Normalized Runtime
2) Performance with different graph partitionings. During the benchmark test, we test 3 different graph partitioning methods with 2, 4, and 8 GPUs, which is shown in Table 7, using gcloud instances with different numbers of V100 GPUs. In order to test the speed of the simulation using different GPUs, we will introduce the notion of ‘Strong Scaling’. Based on Glaser et al. Glaser et al. [2015], in strong scaling, we keep the dataset size (the same demand file in our case) constant but increase the number of processors.
@@
3) Performance with the adjustment of storing vehicles with Device_Vector. Since we performed an improvement of the data structure from array to Device_Vector, we also conduct a test to compare the performance of the simulation time, and the result is shown below in Table 8. We use a computer with two NVIDIA A100 40GB GPUs to run the test. The result shows that simulation time for storing vehicles with an array is far slower than strong vehicles with a Device_Vector. It is mainly because Device_Vectors usually provide dynamic memory allocation and can be scaled at runtime as needed. Arrays, on the other hand, typically require their size to be determined at compile time. In multi-GPU scenarios, dynamic allocation may allow for more efficient memory usage and better memory allocation policies, reducing memory transfers across GPUs. Using an array also requires large host-to-gpu data transfers rather than Device_Vector storage.
@@
4) Roofline Model Result. Figure 13 shows the Instruction Roofline applied to NVIDIA’s A100 for our simulator analysis. Blue dots represent non-predicated instructions for each level of the memory hierarchy, highlight the limited data locality, and speedup from sorting. Gold dots are non-predicated loads per global memory access and highlight the loss in near-unit stride access from sorting. The dotted line represents the total (including predicated) instruction throughput. Proximity of blue dots to the highlighted line quantifies the impact of predication. The A100’s architecture, comprising 108 Streaming Multiprocessors (SMs) with four sub-partitions each, integrates a single warp scheduler capable of dispatching one instruction per cycle within each sub-partition. Consequently, the theoretical maximum instruction throughput on a warp basis is calculated as 108×4×1×1.41 GHz, amounting to 609.12 GIPS. We leverage Yang et al.’s methodology Yang et al. [2019] for measuring GPU memory bandwidths but rescale into gigasectors per second (GSECT/s) based on the sector size. We record the number of instructions executed on the thread level, normalized to warp-level by dividing by 32. This instruction count is then divided by the corresponding sector count in data movement to determine the instruction intensity. Performance metrics are calculated by dividing this instruction count by the kernel execution time. The resulting blue markers determined by these calculations fall beneath the sloping roofline, illustrating that the program is memory-bound with room to boost instruction intensity for better performance. Moreover, the disparity between the L1 marker and the L2 and HBM markers underscores efficient L1 data reuse, but not the L2.

### v2

@@ 3.2.1 GPU Kernel Design for Vehicle and Traffic Management
State: $\mathcal{R}$ (Road Status), $t_{depart}$ (Departure Time of current vehicle), $\mathcal{V}$ (Vehicle Type), $e_{current}$ (Current Edge), $v$ (Speed), $\mathcal{I}$ (Intersection Status)
@@
5:         Start moving, we can write vehicle status on $\mathcal{R}$ and update $\mathcal{P}$ accordingly.
6:         if Moving on a New Edge then
@@
10:              $d_{front}\leftarrow 2\cdot\Delta t\cdot v$ (because maximum distance a vehicle can move within $\Delta t$ is $v\Delta t+\frac{1}{2}a\Delta t^{2}=v\Delta t+\frac{1}{2}v\Delta t=\frac{3}{2}v\Delta t$, we check more than the maximum distance to avoid collision)
@@
18:              Evaluate lane change with Equation (Lane Change)
@@
20:                  if Gap acceptance check with Equation (Gap Acceptancethen
@@
22:return $\mathcal{R}$ , $\mathcal{P},v,a$
In detail, as shown in the Vehicle Propagation Algorithm 1, the $\mathcal{R}$ Road Status array represents the occupancy status or speed of vehicles on the road, while $\mathcal{I}$ (Intersection Status) denotes the graph nodes. The process begins with each thread checking the road status and vehicle’s departure time to decide whether to depart or wait. Once the vehicle is active, the thread orchestrates the vehicle’s progression, revising its position and velocity in response to the immediate road conditions, such as the presence and distance of preceding vehicles. Additionally, the thread evaluates the vehicle’s interaction with intersections and potential lane changes.
@@ 3.3 Design of multiple GPU-based simulation framework
The propagation of vehicles and the complete road network across multiple GPUs calls for inter-GPU communications to synchronize state and share data. When a vehicle is not located in the ghost zone, the simulator checks whether its next node belongs to the ghost zone. If not, the simulation continues to a single-GPU scenario, eliminating the need for cross-GPU computations. Conversely, if the vehicle is about to enter the ghost zone, it is duplicated in the ghost zone of the destination GPU. If a vehicle is already in a ghost zone, it is assured that its next node is out of the ghost zone which means after the current edge traversal, the vehicle will be on the next GPU and leave the original GPU. At this juncture, it is determined whether the node resides within the current GPU’s domain. If it does, the situation reverts to a single-GPU model as described in Session 3.2 again. If the node is outside the current GPU, the vehicle is marked for removal, and subsequently eliminated post the completion of the simulation timestep. The whole procedure is summarized in Figure 4, intersections within the road network are represented by circles, and we use an adjacency list to illustrate the directed relationships between these intersections. The entries in the list link to the arrays that define road segments. For example, a 4-lane 8-meter road segment is depicted as a 1D matrix with a dimension of 1 x 32 as shown in Fig. 4.1, showing how real road segments, represented by lines in Fig. 4, are systematically translated into 1D matrices to sequentially organize lanes. In the process of selecting ghost zones, we replicate the entire edge (represented as a 1D array) across both GPUs and the vehicles on the ghost zone will be calculated simultaneously within two GPUs. To ensure exclusive byte allocation to each vehicle, we transfer vehicles one at a time from one edge to another if the first-byte memory of the downstream edge is not occupied.
@@ 3.3.1 Graph Partitioning
In large-scale multi-GPU parallel simulations, it is essential to distribute graph data across multiple GPUs to share computational resources. We aim for an efficient allocation of extensive graph data, managing computational resources while minimizing inter-GPU communication. In this part, we first state our problem and propose an approximate optimization formulation of the graph partitioning task. Since it is difficult to solve for our network with $200k$ node size, we introduce two applicable graph partitioning strategies to solve the problem. The choice of two graph partitioning methods is closely related to the available computational resources of the single GPU, and the inherent approximation idea of the two methods comes from different perspectives. The framework is summarized as shown in Figure 5, which illustrates that when we have insufficient computational resources, which means that we are creating many more threads than CUDA cores, it leads to a computationally bound scenario. We propose using balanced partitioning to allocate computation evenly across GPUs. In contrast, with ample computational power, the challenge shifts to minimizing communication overhead, necessitating unbalanced partitioning. Our framework consists of three parts: identifying the difficulties in the partition problem, simplifying it to an optimization formulation, and using efficient algorithms to solve the approximation problem. The left side of Fig. 5 is about modeling the problem, the right side is about how to formulate the problem and solve it, and the middle part is about the detailed components of our solutions.
@@
    Outlier Detection: For those nodes never visited in the past time period, we allocate them to the nearest subgraph.
We present an example of partition results for two distinct time periods, showing both balanced and unbalanced graph partitionings in Figures 6 and 7, respectively. These partitions are aligned with the dynamics of the real world traffic flow. Specifically, areas such as the Bay Bridge and Treasure Island are unified within a single cluster due to the incessant flow of traffic in both balanced and unbalanced partition cases because of their heavy traffic throughout the day, underscoring the impracticality of division. In the balanced partition scenario, the map is divided into northern and southern segments by the Bay Bridge area, delineating a clear geographic split. On the other hand, the unbalanced partition strategy, particularly with 4 and 8 clusters, mirrors the Bay Area’s administrative boundaries more closely. For the partition featuring four clusters, the blue, green, and orange parts represent North Bay, East Bay, and South Bay, respectively, while the red segment encompasses the San Francisco and Peninsula areas, along with portions of North and East Bay, interconnected by the Golden Gate Bridge and the Bay Bridge. In the scenario with eight clusters, each segment closely approximates the distinct counties within the Bay Area, with the exception that San Francisco and Marin counties are in one segment.
@@ 3.3.2 Inter-GPU Communication
The additional cost incurred in a multi-GPU 8 setup, compared to a single-GPU system, arises predominantly from the need for communication between GPUs. We facilitate communication between multiple GPUs to handle two types of data: vehicles and road networks within ghost zones. Vehicle data transmission occurs when a vehicle enters a ghost zone, while road networks data communication is necessary to maintain consistency across different GPUs when a vehicle moves within a ghost zone. While the use of arrays to store vehicle data, managed via cudaFree and cudaMalloc, presented an intuitive approach, it revealed significant drawbacks, especially in handling variable data sizes. As illustrated in Scenario 1 of Figure 9, allocating excessive memory leads to inefficiency, whereas insufficient allocation, shown in Scenario 2, poses the risk of array index out of bounds. Moreover, this method was hampered by the time-consuming processes involved in frequent memory allocations and deallocations. To efficiently manage the variable-length vehicle data, which needs to be transferred between GPUs, we employ a more effective approach, utilizing device_vector from the Thrust library, the CUDA C++ template library. Thrust’s device_vector is designed for GPU contexts and allows dynamic resizing of its contained elements. This is achieved through efficient memory management strategies, which minimize the overhead of reallocating memory when the device_vector size changes.
@@
Building on the detailed methodologies for vehicle replication and deletion, and road network data consistency, we’ve established a robust system for managing GPU communication. The following section demonstrates the relatively minimal time impact of communication operations compared to read and write processes from one GPU to its local memory, underscoring the practicality and efficiency of our multi-GPU approach. For read and write operations, the cost was assessed by reading and writing large arrays, where each thread manipulates a single array value. In contrast, communication operations were evaluated by replicating a long array from one GPU to another, using both Device-Host-Device (D-H-D) and Peer-to-Peer (P2P) methods. P2P, especially via NVLink, is claimed faster than D-H-D, though its availability depends on the machine’s hardware architecture. Experimental results on V100 GPUs equipped with NVLink revealed that the time cost for P2P communication is approximately threefold that of read-write operations, shown in Figure 10, indicating a reasonably efficient system. In the upcoming performance section, we will elaborate using the example of San Francisco Bay Area, how graph partitioning techniques and other strategies effectively minimize communication overhead to a mere fraction (about 1‰) of the total computational load. In summary, while inter-GPU communication does introduce additional time costs, this is relatively minimal compared to the extensive read-and-write operations, thereby justifying the use of multiple GPUs for their significant performance benefits.
@@
Figure 11: The Demonstration of Vehicle Replication and Deletion
@@ 4.1 Experimental Data Source
Our analysis focused on a comprehensive dataset encompassing over 23.5 million recorded trips, which is a typical weekday with no significant events or seasonal impactsChan et al. [2023]. Each entry in this dataset was categorized by origin, destination, mode of transportation, and several other critical attributes. A significant portion of our research was devoted to examining car trips, specifically categorized as Single-Occupancy Vehicle (SOV), High Occupancy Vehicle with two occupants (HOV 2), and High Occupancy Vehicle with three or more occupants (HOV 3+). Upon applying these criteria, our scope narrowed down to 17.8 million car trips, constituting 75.7% of the total dataset. This subset provided a substantial basis for analyzing vehicular movement patterns within the region, offering insights into the predominant vechicle propagation and its implications for urban planning and traffic management.
@@ 4.2 Experimental Results
1) Performance with different graph partitionings. During the benchmark test, we test 3 different graph partitioning methods with 2, 4, and 8 GPUs, which is shown in Table 7, using gcloud instances with different numbers of V100 GPUs. In order to test the speed of the simulation using different GPUs, we will introduce the notion of ‘Strong Scaling’. Based on Glaser et al. Glaser et al. [2015], in strong scaling, we keep the dataset size (the same demand file in our case) constant but increase the number of processors.
@@
2) Performance with the adjustment of storing vehicles with Device_Vector. Since we performed an improvement of the data structure from array to Device_Vector, we also conduct a test to compare the performance of the simulation time, and the result is shown below in Table 5. We use a computer with two NVIDIA A100 40GB GPUs to run the test. The result shows that simulation time for storing vehicles with an array is far slower than storing vehicles with a Device_Vector. It is mainly because Device_Vectors usually provide dynamic memory allocation and can be scaled at runtime as needed. Arrays, on the other hand, typically require their size to be determined at compile time. In multi-GPU scenarios, dynamic allocation may allow for more efficient memory usage and better memory allocation policies, reducing memory transfers across GPUs. Using an array also requires large host-to-gpu data transfers rather than Device_Vector storage.
@@
3) Roofline Model Result. Figure 12 shows the Instruction Roofline applied to NVIDIA’s A100 for our simulator analysis. Blue dots represent non-predicated instructions for each level of the memory hierarchy, highlight the limited data locality, and speedup from sorting. Gold dots are non-predicated loads per global memory access and highlight the loss in near-unit stride access from sorting. The dotted line represents the total (including predicated) instruction throughput. Proximity of blue dots to the highlighted line quantifies the impact of predication. The A100’s architecture, comprising 108 Streaming Multiprocessors (SMs) with four sub-partitions each, integrates a single warp scheduler capable of dispatching one instruction per cycle within each sub-partition. Consequently, the theoretical maximum instruction throughput on a warp basis is calculated as 108×4×1×1.41 GHz, amounting to 609.12 GIPS. We leverage Yang et al.’s methodology Yang et al. [2019] for measuring GPU memory bandwidths but rescale into gigasectors per second (GSECT/s) based on the sector size. We record the number of instructions executed on the thread level, normalized to warp-level by dividing by 32. This instruction count is then divided by the corresponding sector count in data movement to determine the instruction intensity. Performance metrics are calculated by dividing this instruction count by the kernel execution time. The resulting blue markers determined by these calculations fall beneath the sloping roofline, illustrating that the program is memory-bound with room to boost instruction intensity for better performance. Moreover, the disparity between the L1 marker and the L2 and HBM markers underscores efficient L1 data reuse, but not the L2.
Before we dive into Performance Analysis with Roofline Model, we will first introduce the concept of thread predication. Thread predication in GPUs is a mechanism designed to handle conditional branches, allowing the SIMD architecture to continue operating even when threads within a warp follow different execution paths. This section provides a detailed explanation of how thread predication works and its impact on performance.
Concept of Thread Predication: In GPU architecture, a warp is a group of threads that execute instructions simultaneously. When a conditional branch (e.g., an if-else statement) occurs, different threads in the same warp might need to take different execution paths based on their individual conditions. Since a warp executes the same instruction at the same time across all its threads, this presents a challenge.
Mechanism of Thread Predication: When the warp encounters a conditional branch, it must execute both paths of the branch to ensure all threads complete their respective instructions. The GPU uses predication to manage this. It deactivates (predicates off) the threads that do not meet the condition for the current branch, while the active threads execute the instructions for that branch. Once this is done, the GPU switches to the other branch, activating the previously deactivated threads and deactivating those that have already executed their instructions. Consequently, the warp processes each branch sequentially, with only a subset of threads active at any one time.
For instance, consider the following example:
if (condition) {
    // Branch A: executed by threads where ’condition’ is true
    doSomething();
} else {
    // Branch B: executed by threads where ’condition’ is false
    doSomethingElse();
If the condition is true for half of the threads in a warp, those threads will execute doSomething() while the other half are deactivated. Then, the roles are reversed, and the second half executes doSomethingElse().
Impact on Performance: The impact of thread predication on performance is significant. Since only a portion of the warp’s threads are active at any given time, the GPU’s parallel processing capability is underutilized. For instance, if only half the threads are active, the warp operates at 50% efficiency. Each branch must be executed separately for the different sets of active threads, effectively doubling the instruction count for divergent branches. This sequential processing of branches increases the overall execution time, leading to a performance bottleneck when there is significant divergence.
Performance Analysis with Roofline Model: Thread predication deactivates threads not following a branch, impacting performance by limiting the active thread count in a warp. Since a warp processes a single instruction collectively, the fraction of active threads directly influences efficiency.